# Random dynamics and invariant measures for a class of non-Newtonian fluids of differential type on 2D and 3D Poincaré domains

### v1

@@ Abstract.
In this article, we consider a class of incompressible stochastic third-grade fluids (non-Newtonian fluids) equations on two- as well as three-dimensional Poincaré domains $\mathscr{O}$ (which may be bounded or unbounded). Our aims are to study the well-posedness and asymptotic analysis for the solutions of the underlying system. Firstly, we prove that the underlying system defined on $\mathscr{O}$ has a unique weak solution (in the analytic sense) under Dirichlet boundary condition and it also generates random dynamical system $\Psi$. Secondly, we consider the underlying system on bounded domains. Using the compact Sobolev embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$, we prove the existence of a unique random attractor for the underlying system on bounded domains with external forcing in $\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$. Thirdly, we consider the underlying system on unbounded Poincaré domains with external forcing in $\mathbb{L}^{2}(\mathscr{O})$ and show the existence of a unique random attractor. In order to obtain the existence of a unique random attractor on unbounded domains, due to the lack of compact Sobolev embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$, we use the uniform-tail estimates method which helps us to demonstrate the asymptotic compactness of $\Psi$. Note that due to the presence of several nonlinear terms in the underlying system, we are not able to use the energy equality method to obtain the asymptotic compactness of $\Psi$ in unbounded domains, which makes the analysis of this work in unbounded domains more difficult and interesting. Finally, as a consequence of the existence of random attractors, we address the existence of invariant measures for underlying system. To the best of authors’ knowledge, this is the first work which consider the 2D as well as 3D incompressible stochastic third-grade fluids equations and establish the existence of random attractor in bounded as well as unbounded domains. In addition, this is the first work which address the existence of invariant measures for underlying system on unbounded domains.
@@ 1.1. The underlying system
$\partial_{t}(z(\boldsymbol{v}))-\nu\Delta\boldsymbol{v}+(\boldsymbol{v}\cdot\nabla)z(\boldsymbol{v})&+\displaystyle\sum_{j=1}^{d}z(\boldsymbol{v})^{j}\nabla v^{j}-(\alpha_{1}+\alpha_{2})\text{div}((\mathrm{A}(\boldsymbol{v}))^{2})\\ -\beta\text{div}[\mathrm{Tr}(\mathrm{A}(\boldsymbol{v})\mathrm{A}(\boldsymbol{v})^{T})\mathrm{A}(\boldsymbol{v})]&=-\nabla\mathbf{P}+\boldsymbol{f},\\[5.69046pt] \text{div}\;\boldsymbol{v}&=0,\\[5.69046pt] z(\boldsymbol{v})&:=\boldsymbol{v}-\alpha_{1}\Delta\boldsymbol{v},\\[5.69046pt] \mathrm{A}(\boldsymbol{v})&:=\nabla\boldsymbol{v}+(\nabla\boldsymbol{v})^{T},\end{aligned}\right.$ (1.3)
@@
Our aim in this work is to study the following stochastic version of system (1.3) with a subset of physical conditions (1.2), namely
@@
$\mathrm{d}\boldsymbol{v}&=[\boldsymbol{f}-\nabla\textbf{P}+\nu\Delta\boldsymbol{v}-(\boldsymbol{v}\cdot\nabla)\boldsymbol{v}+\alpha\text{div}((\mathrm{A}(\boldsymbol{v}))^{2})+\beta\text{div}(|\mathrm{A}(\boldsymbol{v})|^{2}\mathrm{A}(\boldsymbol{v}))]\mathrm{d}t&&\\ &\quad+\mathrm{d}\mathrm{W}&&\text{in }\mathscr{O}\times(0,\infty),\\ \text{div}\;\boldsymbol{v}&=0\quad&&\text{in }\mathscr{O}\times[0,\infty),\\ \boldsymbol{v}&=\boldsymbol{0}&&\text{on }\partial\mathscr{O}\times[0,\infty),\\ \boldsymbol{v}(x,0)&=\boldsymbol{x}\quad&&\text{in }\mathscr{O},\end{aligned}\right.$ (1.5)
@@ 1.3. Scope of the article, approaches and assumptions
Broadly, this article is divided into three main parts. Our first aim is to establish the existence and uniqueness of weak solutions (in the analytic sense) of the stochastic third-grade fluids equations (1.5). We begin by defining an Ornstein-Uhlenbeck process which takes values in $\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})$. In order to define a suitable Ornstein-Uhlenbeck process, one of the technical difficulties related to the noise is resolved by using the corresponding RKHS. In particular, motivated from [11, 12, 13] for 2D NSE on unbounded Poincaré domains, we assume that the RKHS $\mathrm{K}$ satisfies the following hypothesis:
@@
[…]
@@
[…]
The condition (1.7) also says that the mapping $\mathscr{A}^{-\delta-1-\frac{r}{2}}:\mathrm{K}\to\mathbb{H}$ is Hilbert-Schmidt. If $\mathscr{O}$ is a bounded domain, then $\mathscr{A}^{-s}:\mathbb{H}\to\mathbb{H}$ is Hilbert-Schmidt if and only if $\sum_{j=1}^{\infty}\lambda_{j}^{-2s}<+\infty,$ where $\mathscr{A}e_{j}=\lambda_{j}e_{j}$, $j\in\mathbb{N}$ and $\{e_{j}\}_{j\in\mathbb{N}}$ is an orthogonal basis of $\mathbb{H}$. In bounded domains, it is well known that $\lambda_{j}\sim j^{\frac{2}{d}},$ for large $j$ (growth of eigenvalues) and hence $\mathscr{A}^{-s}$ is Hilbert-Schmidt if and only if $s>\frac{d}{4}.$ In other words, with $\mathrm{K}=\mathrm{D}(\mathscr{A}^{s+1+\frac{r}{2}}),$ the embedding $\mathrm{K}\hookrightarrow\mathrm{D}(\mathscr{A}^{1+\frac{r}{2}})\hookrightarrow\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})$ is $\gamma$-radonifying if and only if $s>\frac{d}{4}.$ Thus, Hypothesis 1.1 is satisfied for any $\delta>0.$ In fact, the condition (1.6) holds if and only if the operator $\mathscr{A}^{-(s+1+\frac{r}{2}+\delta)}:\mathbb{H}\to\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})$ is $\gamma$-radonifying.
@@
Making use of the Ornstein-Uhlenbeck process, we define a Doss-Sussman transformation [29, 59] (see (3.17) below) and obtain an equivalent pathwise deterministic system (see system (3.18) below) to system (1.5). Next, we show the existence of a unique weak solution to the equivalent pathwise deterministic system (3.18) via a monotonicity argument and a Minty-Browder technique. The following monotonicity result of the operator $\mathscr{G}(\cdot):=\nu\mathscr{A}\cdot+\mathscr{B}(\cdot+\boldsymbol{z})+\alpha\mathscr{J}(\cdot+\boldsymbol{z})+\beta\mathscr{K}(\cdot+\boldsymbol{z})$ plays a crucial role:
$\mathscr{G}(\boldsymbol{y}_{1})-\mathscr{G}(\boldsymbol{y}_{2}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>+\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})\|_{{4}}^{2}\|\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\|_{2}^{2}$
@@
Secondly, we are interested in the large-time behavior of solutions of system (1.5), that is, the existence of random attractors for system (1.5). Similar to the case of NSE, it is a challenging problem to study the long-time behavior of third-grade fluids equations on the whole space $\mathbb{R}^{d}$. Nevertheless, one can try these kinds of analysis on Poincaré domains which may be bounded or unbounded, see for instance [12, 13, 43, 67], etc. By a Poincaré domain, we mean a domain in which the Poincaré inequality is satisfied. More precisely, we consider the following hypothesis on the domain $\mathscr{O}$:
@@
Let $\mathscr{O}$ be an open, connected and unbounded subset of $\mathbb{R}^{3}$, the boundary of which is uniformly of class $\mathrm{C}^{3}$ (see [35]). We assume that, there exists a positive constant $\lambda$ such that the following Poincaré inequality is satisfied:
@@
A typical example of a unbounded Poincaré domain in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ are $\mathscr{O}=\mathbb{R}\times(-L,L)$ and $\mathscr{O}=\mathbb{R}^{2}\times(-L,L)$ with $L>0$, respectively, see [62, p.306] and [57, p.117].
@@
In this work, we apply the abstract theory established in [13] to demonstrate the existence of random attractors for system (1.5). In order to use the abstract result, we first show that system (1.5) generates random dynamical system $\Psi$ (see Section 3). Next, we need to show that $\Psi$ has an random absorbing set and it is asymptotically compact. We show the existence of random absorbing set using the uniform energy estimates. On the other hand, on bounded domains, we use the compact Sobolev embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$ to obtain asymptotic compactness of $\Psi$. In unbounded domains, due to lack of compact Sobolev embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$, the commonly use methods in literature are the following:
@@
Note that unlike the Navier-Stokes equations (see [12, 13]), convective Brinkman-Forchheimer equations (see [38, 41], etc.), etc. we are not able to show that the solutions of third-grade fluids equations (1.5) satisfy the weak continuity property with respect to the initial data. Due to this reason, we are not able to use the energy equality method to obtain the asymptotic compactness of $\Psi$ on unbounded domains. Therefore, for unbounded domains, we explore the uniform-tail estimate method to prove the asymptotic compactness of $\Psi$ (Theorem 5.5).
@@
In the context of attractor when $\mathscr{O}$ is a bounded domains, we have considered $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$. But for unbounded domain case, we need to restrict the forcing term to be in a regular space, that is, $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$. If one proves the weak continuity of the solutions with respect to the initial data of the third-grade fluids equation, it will be possible to work with the same regularity as in bounded domains via energy equality method (see the works [13, 38], etc. for Navier-Stokes equations and related models). Unfortunately, we are not able to prove the weak continuity of the solutions of third-grade fluids equations with respect to the initial data, so the method of energy equality is not applied.
Essentially unlike the parabolic or hyperbolic equations as considered in [18, 48, 64, 68], etc. the fluids equations like (1.5) contain the pressure term P. When we derive these uniform tail-estimates, the pressure term P can not be simply treated by the divergence theorem (does not vanish). However, by taking the divergence in (1.5) and using the incompressibility condition, we get the rigorous expression of the pressure term
P $=(-\Delta)^{-1}\nabla\cdot\left[\nabla\cdot(\boldsymbol{v}\otimes\boldsymbol{v})\right]+\alpha(-\Delta)^{-1}\nabla\cdot\left[\nabla\cdot\left\{(\mathrm{A}(\boldsymbol{v}))^{2}\right\}\right]$
$\quad+\beta(-\Delta)^{-1}\nabla\cdot\left[\nabla\cdot\left\{|\mathrm{A}(\boldsymbol{v})|^{2}\mathrm{A}(\boldsymbol{v})\right\}\right]-(-\Delta)^{-1}\left[\nabla\cdot\boldsymbol{f}\right].$
Then it is possible to derive uniform-tail estimates by estimating all nonlinear terms in system (1.5) and pressure P carefully, see Lemma 5.4 below. As a result of uniform-tail estimates, the asymptotic compactness of $\Psi$ on unbounded domains follows (see Theorem 5.5).
@@
    there exists a unique weak solution (in the analytic sense) for the stochastic third-grade fluids equations (1.5) with external forcing $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$,
@@
    on bounded domains, there exists a unique random attractor for the stochastic third-grade fluids equations (1.5) with external forcing $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$,
@@
    on unbounded domains, there exists a unique random attractor for the stochastic third-grade fluids equations (1.5) with external forcing $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$,
@@
    on bounded and unbounded domains, there exists an invariant measure for the stochastic third-grade fluids equations (1.5) with external forcing $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})+\mathbb{W}^{-1,\frac{4}{3}}(\mathscr{O})$ and $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$, respectively.
We remark that to the best of authors’ knowledge, the well-posedness and the large-time behavior of the solutions to the stochastic third-grade fluids equations have not been investigated in bounded/unbounded domains till the date. In addition, the existence of invariant measures for the stochastic third-grade fluids equations on unbounded domains is being studied here for the first time. Unfortunately, we are unable to demonstrate the uniqueness of invariant measures, which is a subject we plan to tackle in the future.

### v2

@@ Abstract.
In this article, we consider a class of incompressible stochastic third-grade fluids (non-Newtonian fluids) equations on two- as well as three-dimensional Poincaré domains $\mathscr{O}$ (which may be bounded or unbounded). Our aims are to study the well-posedness and asymptotic analysis for the solutions of the underlying system. Firstly, we prove that the underlying system defined on $\mathscr{O}$ has a unique weak solution (in the analytic sense) under Dirichlet boundary condition and it also generates random dynamical system $\Psi$. Secondly, we consider the underlying system on bounded domains. Using the compact Sobolev embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$, we prove the existence of a unique random attractor for the underlying system on bounded domains with external forcing in $\mathbb{H}^{-1}(\mathscr{O})$. Thirdly, we consider the underlying system on unbounded Poincaré domains with external forcing in $\mathbb{L}^{2}(\mathscr{O})$ and show the existence of a unique random attractor. In order to obtain the existence of a unique random attractor on unbounded domains, due to the lack of compact Sobolev embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$, we use the uniform-tail estimates method which helps us to demonstrate the asymptotic compactness of $\Psi$. Note that due to the presence of several nonlinear terms in the underlying system, we are not able to use the energy equality method to obtain the asymptotic compactness of $\Psi$ in unbounded domains, which makes the analysis of this work in unbounded domains more difficult and interesting. Finally, as a consequence of the existence of random attractors, we address the existence of invariant measures for underlying system. To the best of authors’ knowledge, this is the first work which consider a class of the 2D as well as 3D incompressible stochastic third-grade fluids equations and establish the existence of random attractor in bounded as well as unbounded domains. In addition, this is the first work which address the existence of invariant measures for underlying system on unbounded domains.
@@ 1.1. The underlying system
$\left\{\begin{aligned} \partial_{t}(z(\boldsymbol{v}))-\nu\Delta\boldsymbol{v}+(\boldsymbol{v}\cdot\nabla)z(\boldsymbol{v})&+\displaystyle\sum_{j=1}^{d}z(\boldsymbol{v})^{j}\nabla v^{j}-(\alpha_{1}+\alpha_{2})\text{div}((\mathrm{A}(\boldsymbol{v}))^{2})\\ -\beta\text{div}[\mathrm{Tr}(\mathrm{A}(\boldsymbol{v})\mathrm{A}(\boldsymbol{v})^{T})\mathrm{A}(\boldsymbol{v})]&=-\nabla\mathbf{P}+\boldsymbol{f},\\[5.69046pt] \text{div}\;\boldsymbol{v}&=0,\\[5.69046pt] z(\boldsymbol{v})&:=\boldsymbol{v}-\alpha_{1}\Delta\boldsymbol{v},\\[5.69046pt] \mathrm{A}(\boldsymbol{v})&:=\nabla\boldsymbol{v}+(\nabla\boldsymbol{v})^{T},\end{aligned}\right$. (1.3)
@@
Let $\mathscr{O}$ be an open and connected subset of $\mathbb{R}^{d}$ ($d\in\{2,3\}$). Our aim in this work is to study the following stochastic version of system (1.3) with a subset of physical conditions (1.2), namely
@@
$\left\{\begin{aligned} \mathrm{d}\boldsymbol{v}&=[\boldsymbol{f}-\nabla\textbf{P}+\nu\Delta\boldsymbol{v}-(\boldsymbol{v}\cdot\nabla)\boldsymbol{v}+\alpha\text{div}((\mathrm{A}(\boldsymbol{v}))^{2})+\beta\text{div}(|\mathrm{A}(\boldsymbol{v})|^{2}\mathrm{A}(\boldsymbol{v}))]\mathrm{d}t&&\\ &\quad+\mathrm{d}\mathrm{W}&&\hskip-42.67912pt\text{in }\mathscr{O}\times(0,\infty),\\ \text{div}\;\boldsymbol{v}&=0\quad&&\hskip-42.67912pt\text{in }\mathscr{O}\times[0,\infty),\\ \boldsymbol{v}&=\boldsymbol{0}&&\hskip-42.67912pt\text{on }\partial\mathscr{O}\times[0,\infty),\\ \boldsymbol{v}(x,0)&=\boldsymbol{x}\quad&&\hskip-42.67912pt\text{in }\mathscr{O},\end{aligned}\right$. (1.5)
@@ 1.3. Scope of the article, approaches and assumptions
Broadly, this article is divided into three main parts. Our first aim is to establish the existence and uniqueness of weak solutions (in the analytic sense) of system (1.5). We begin by defining an Ornstein-Uhlenbeck process which takes values in $\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})$. In order to define a suitable Ornstein-Uhlenbeck process, one of the technical difficulties related to the noise is resolved by using the corresponding RKHS. In particular, motivated from [12, 13, 10] for 2D NSE on unbounded Poincaré domains, we assume that the RKHS $\mathrm{K}$ satisfies the following hypothesis:
@@
$\mathscr{A}^{-\delta}:\mathrm{K}\rightarrow\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})\ \text{ is }\ \gamma\text{-radonifying.}$ (1.6)
@@
$\mathscr{A}^{-\delta-1-\frac{r}{2}}:\mathrm{K}\rightarrow\mathbb{H}\ \text{ is }\ \gamma\text{-radonifying.}$ (1.7)
The condition (1.7) also says that the mapping $\mathscr{A}^{-\delta-1-\frac{r}{2}}:\mathrm{K}\rightarrow\mathbb{H}$ is Hilbert-Schmidt. If $\mathscr{O}$ is a bounded domain, then $\mathscr{A}^{-s}:\mathbb{H}\rightarrow\mathbb{H}$ is Hilbert-Schmidt if and only if $\sum_{j=1}^{\infty}\lambda_{j}^{-2s}<+\infty,$ where $\mathscr{A}e_{j}=\lambda_{j}e_{j}$, $j\in\mathbb{N}$ and $\{e_{j}\}_{j\in\mathbb{N}}$ is an orthogonal basis of $\mathbb{H}$. In bounded domains, it is well known that $\lambda_{j}\sim j^{\frac{2}{d}},$ for large $j$ (growth of eigenvalues) and hence $\mathscr{A}^{-s}$ is Hilbert-Schmidt if and only if $s>\frac{d}{4}.$ In other words, with $\mathrm{K}=\mathrm{D}(\mathscr{A}^{s+1+\frac{r}{2}}),$ the embedding $\mathrm{K}\hookrightarrow\mathrm{D}(\mathscr{A}^{1+\frac{r}{2}})\hookrightarrow\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})$ is $\gamma$-radonifying if and only if $s>\frac{d}{4}.$ Thus, Hypothesis 1.1 is satisfied for any $\delta>0.$ In fact, the condition (1.6) holds if and only if the operator $\mathscr{A}^{-(s+1+\frac{r}{2}+\delta)}:\mathbb{H}\rightarrow\mathbb{H}\cap{\mathbb{W}}^{1,4}(\mathscr{O})$ is $\gamma$-radonifying.
@@
Making use of the Ornstein-Uhlenbeck process, we define a Doss-Sussmann transformation [30, 67] (see (3.17) below) and obtain an equivalent pathwise deterministic system (see system (3.18) below) to system (1.5). Next, we show the existence of a unique weak solution to the equivalent pathwise deterministic system (3.18) via a monotonicity argument and a Minty-Browder technique. The following monotonicity result of the operator $\mathscr{G}(\cdot):=\nu\mathscr{A}\cdot+\mathscr{B}(\cdot+\boldsymbol{z})+\alpha\mathscr{J}(\cdot+\boldsymbol{z})+\beta\mathscr{K}(\cdot+\boldsymbol{z})$ plays a crucial role:
$\left<\mathscr{G}(\boldsymbol{y}_{1})-\mathscr{G}(\boldsymbol{y}_{2}),\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\right>+\frac{(C_{S,3}C_{K})^{2}}{\nu\varepsilon_{0}}\|\mathrm{A}(\boldsymbol{y}_{2}+\boldsymbol{z})\|_{{4}}^{2}\|\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\|_{2}^{2}$
@@
Secondly, we are interested in the large-time behavior of solutions of system (1.5), that is, the existence of random attractors for system (1.5). Similar to the case of NSE, it is a challenging problem to study the long-time behavior of system (1.5) on the whole space $\mathbb{R}^{d}$. Nevertheless, one can try these kinds of analysis on Poincaré domains which may be bounded or unbounded, see for instance [13, 10, 47, 77], etc. By a Poincaré domain, we mean a domain in which the Poincaré inequality is satisfied. More precisely, we consider the following hypothesis on the domain $\mathscr{O}$:
@@
Let $\mathscr{O}$ be an open and connected subset (may be bounded or unbounded) of $\mathbb{R}^{d}$, $d\in\{2,3\}$, the boundary of which is uniformly of class $\mathrm{C}^{1}$ (see [34]). We assume that, there exists a positive constant $\lambda$ such that the following Poincaré inequality is satisfied:
@@
A typical example of a unbounded Poincaré domain in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ are $\mathscr{O}=\mathbb{R}\times(-L,L)$ and $\mathscr{O}=\mathbb{R}^{2}\times(-L,L)$ with $L>0$, respectively, see [71, p.306] and [63, p.117].
@@
In this work, we apply the abstract theory established in [10] to demonstrate the existence of random attractors for system (1.5). In order to use the abstract result, we first show that system (1.5) generates random dynamical system $\Psi$ (see Section 3). Next, we need to show that $\Psi$ has a random absorbing set and it is asymptotically compact. We show the existence of random absorbing set using the uniform energy estimates. On the other hand, on bounded domains, we use the compact Sobolev embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$ to obtain asymptotic compactness of $\Psi$. In unbounded domains, due to lack of compact Sobolev embedding $\mathbb{H}^{1}(\mathscr{O})\hookrightarrow\mathbb{L}^{2}(\mathscr{O})$, the commonly use methods in literature are the following:
@@
Note that unlike the Navier-Stokes equations (see [13, 10]), convective Brinkman-Forchheimer equations (see [42, 45], etc.), etc. we are not able to show that the solutions of system (1.5) satisfy the weak continuity property with respect to the initial data. Due to this reason, we are not able to use the energy equality method to obtain the asymptotic compactness of $\Psi$ on unbounded domains. Therefore, for unbounded domains, we explore the uniform-tail estimate method to prove the asymptotic compactness of $\Psi$ (Theorem 5.5).
@@
In the context of attractor when $\mathscr{O}$ is a bounded domains, we have considered $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})$. But for unbounded domain case, we need to restrict the forcing term to be in a regular space, that is, $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$. If one proves the weak continuity of the solutions with respect to the initial data of system (1.5), it will be possible to work with the same regularity as in bounded domains via energy equality method (see the works [10, 42], etc. for Navier-Stokes equations and related models). Unfortunately, we are not able to prove the weak continuity of the solutions of system (1.5) with respect to the initial data, so the method of energy equality is not applied.
Remark 1.5.
Comparing with the Navier-Stokes equations, system (1.5) has stronger nonlinear terms, nevertheless due to its symmetry, the nonlinear term $\text{div}(|\mathrm{A}(\boldsymbol{v})|^{2}\mathrm{A}(\boldsymbol{v}))$ has an important regularizing effect, ensuring uniqueness even in 3D. At first glance, one might expect that system (1.5) would also possess the weak continuity property, however this property is not easy to show. To clarify this issue, let $\boldsymbol{v}_{n}(\cdot)$ denote the unique solution of system (1.5) corresponding to the initial data $\boldsymbol{v}_{0,n}$, for $n\in\mathbb{N}$. By definition, weak continuity requires that if $\boldsymbol{v}_{0,n}$ converges weakly to $\boldsymbol{v}_{0}$ in $\mathbb{H}$, then $\boldsymbol{v}_{n}(\cdot)$ converges weakly to $\boldsymbol{v}(\cdot)$ in $\mathbb{H}$, where $\boldsymbol{v}(\cdot)$ is the unique solution of system (1.5) associated with initial data $\boldsymbol{v}_{0}$. However, due to the gradient-type nonlinearity, it is necessary to apply the monotonicity method to pass to the limit $n\rightarrow\infty$ in (1.5), which requires strong convergence of the initial data. The monotonicity method does not appear suitable for establishing weak continuity of solutions to system (1.5). In the Navier-Stokes equations case, the limit can be performed without relying on monotonicity method.
Essentially unlike the parabolic or hyperbolic equations as considered in [19, 54, 73, 78], etc. the fluids equations like (1.5) contain the pressure term P. When we derive these uniform tail-estimates, the pressure term P can not be simply treated by the divergence theorem (does not vanish). We derive uniform-tail estimates by estimating all nonlinear terms in system (1.5) and pressure P carefully, see Lemma 5.4 below. As a result of uniform-tail estimates, the asymptotic compactness of $\Psi$ on unbounded domains follows (see Theorem 5.5).
@@
    there exists a unique weak solution (in the analytic sense) for system (1.5) with external forcing $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})$,
@@
    on bounded domains, there exists a unique random attractor for system (1.5) with external forcing $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})$,
@@
    on unbounded domains, there exists a unique random attractor for system (1.5) with external forcing $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$,
@@
    on bounded and unbounded domains, there exists an invariant measure for system (1.5) with external forcing $\boldsymbol{f}\in\mathbb{H}^{-1}(\mathscr{O})$ and $\boldsymbol{f}\in\mathbb{L}^{2}(\mathscr{O})$, respectively.
We remark that to the best of authors’ knowledge, the large-time behavior of the solutions to the subclass of stochastic third-grade fluids equations (1.3) under analysis (see system (1.5)) has not been investigated in bounded/unbounded domains till the date. In addition, the existence of invariant measures for this subclass of stochastic third-grade fluids equations (1.3) on unbounded domains is being studied here for the first time. It is important to mention that in 2D and 3D bounded domains, this class of fluids has been considered in the article [69], where the authors proved well-posedness and existence of an ergodic invariant measure for multiplicative Lipschitz noise but not discussed the random dynamics of solutions. We note that, in the case of bounded domains, a direct consequence of the existence of a random attractor yields a particular case of the result established in [69]. Unfortunately, we are unable to demonstrate the uniqueness of invariant measures, which is a subject we plan to tackle in the future.
Remark 1.6.
It is worth mentioning that the notion of pathwise random attractors was first introduced in [28, 36], and subsequently, several authors employed this framework to establish the existence of random attractors for various SPDEs. However, it has been consistently observed that the existing works on pathwise random attractors confine the noise to either additive form, linear multiplicative form, or to coefficients with specific structural properties such as antisymmetry (see [49, 76]). This restriction arises because, in such settings, a random dynamical system (or random cocycle) can be defined via random transformations, most notably the Doss-Sussmann transformation [67]. To the best of our knowledge, no results are available in the literature regarding the existence of pathwise random attractors for SPDEs with general Lipschitz nonlinear diffusion coefficients, an issue that poses a significant challenge. To address this difficulty, the concepts of mean-square and weak mean-square random attractors were introduced in [49], although these approaches remain quite restrictive in practice (see [76] for further discussion). More recently, to handle Lipschitz nonlinear diffusion terms, the concept of weak pullback mean random attractors in spaces of Bochner integrable functions was developed in [76], where the long-term dynamics of stochastic reaction-diffusion equations with both nonlinear drift and nonlinear diffusion were studied. The existence of pathwise random attractors for noise beyond the additive case, as well as the existence of weak pullback mean random attractors in the presence of nonlinear noise coefficients, remains an interesting open problem for the underlying system. These issues will be addressed in future work.